Circular No. 8207

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ig IN UGC 2971 Further to IAUC 8206, J. Graham and W. Li report the LOTOSS

discovery of an apparent supernova on unfiltered KAIT images taken on Sept. 22.5 (at mag 18.4) and 23.5 UT (at mag 18.2). SN 2003ig is located at α = 4^{h}09^{m}43^{s}.50, δ = +37^{o}00′45′′.8 (equinox 2000.0), which is 2′′.0 east and 11′′.0 north of the nucleus of UGC 2971. KAIT images taken on Sept. 11.5 (limiting mag ∼ 18) and on Feb. 8.2 (limiting mag ∼ 19.5) showed nothing at this position.

SUPERNOVA 2003ih IN UGC 2836 Further to IAUC 8200, M. Armstrong reports his discovery of an ap-

parent supernova (mag ∼ 18.1) on unfiltered CCD images (limiting mag ∼ 20–21) taken on Sept. 15.04, 21.04, and 23.16 UT; Armstrong’s position for the new object is α = 3^{h}43^{m}57^{s}.72, δ = +39^{o}17′43′′.9 (equinox 2000.0), or 4′′ east and 1′′ north of the center of UGC 2836. Armstrong adds that nothing is present at this location on Palomar Sky Survey images. W. Li notes that KAIT images show a faint new object (mag ∼ 19.7) on Sept. 19.5, located at position end figures 57^{s}.72, 43′′.2 (or 9′′.4 east and 1′′.0 north of the galaxy nucleus). Nothing is visible at this location on KAIT images taken on Feb. 2.1 (limiting mag ∼ 20.5) or Sept. 8.5 (limiting mag ∼ 19.5). SN 2001I also appeared in this galaxy (cf. IAUC 7563).

SUPERNOVA 2003ie IN NGC 4051 S. Benetti, H. Navasardyan, A. Pastorello, V. Botte, S. Ciroi, G. Al-

tavilla, N. Elias-Rosa, M. Riello, M. Turatto, L. Zampieri, and E. Giro, Osservatorio Astronomico di Padova; and E. Cappellaro, Osservatorio Astronomico di Capodimonte, write that inspection of a spectrum (range 370– 750 nm; resolution 2.4 nm), obtained on Sept. 22.78 UT with the Asiago 1.8-m telescope (+ AFOSC), shows that SN 2003ie (IAUC 8205) is a type-II supernova. The spectrum is dominated by a red continuum streaked by unusually strong P-Cyg lines of Fe II at 492.4, 501.8, and 516.9 nm; Sc II at 552.7, 565.8, and 624.5 nm; Na I D; and Ba II at 614.2 nm. The mean expansion velocity deduced from these lines is 5450 ± 200 km/s. Hβ and Hα are also present with P-Cyg profiles and expansion velocities of 5700 and 5500 km/s, respectively. A broad and asymmetric absorption centered at 629.0 nm does not have a clear identification, but it could be due to a blend of several metal (mostly Ti II and Ba II) lines.

2003 September 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green